OpenStudy (anonymous):

What is the common ratio for the sequence 256, 128, 64, 32, ...? 1/4 1/2 2 4 pls help

OpenStudy (anonymous):

I got 1/4

OpenStudy (anonymous):

This is geometric series with first term, a = 256, common ratio, r = 1/2 nth term, an = ar^(n-1) a_11 = 256(1/2)^10 = 2^8 /2^10 = 1/4 = 0.25

OpenStudy (andrewishere):

I think is is 1/2

OpenStudy (anonymous):

I just worked it out. and I got 1/4 or 0.25

OpenStudy (andrewishere):

256*1/2 is 128 128*1/2 is 64 64*1/2 is 32

OpenStudy (andrewishere):

If you do it backwards, you would multiply by 2 so... 32*2=64 64*2=128 128*2=256

OpenStudy (andrewishere):

So it is 1/2

OpenStudy (anonymous):

Ok, I see now. So the answer is 1/2

OpenStudy (andrewishere):

Yes :)

OpenStudy (anonymous):

ayyy thanks

OpenStudy (andrewishere):

Np I hope that makes sense

OpenStudy (anonymous):

i did